# SplatSuRe: Selective Super-Resolution for Multi-view Consistent 3D Gaussian Splatting

### v1

@@ 6 Results
Figures 1 and 5 present representative renderings from our method. Our approach produces consistently sharper reconstructions while preserving smoothness in uniformly textured regions. In contrast, Mip-Splatting yields overly blurred results because it is trained only on LR and its 3D anti-aliasing filter excessively blurs all Gaussians. SRGS recovers higher-frequency content in some areas but remains constrained by the underlying SR images – regions where the SR output is consistent appear sharp, whereas areas with view-inconsistent SR predictions become noticeably blurred or distorted.
@@
Table 1 presents results on Tanks & Temples [Knapitsch2017tandt]. Our method achieves the highest image quality across nearly all metrics, demonstrating both its effectiveness and robustness. The only exception is NIQE, where 3DGS attains a better score. Similarly, SRGS ranks second across most reference and perceptual metrics but lags behind in MUSIQ and NIQE, where 3DGS performs better. This trend arises because the aliased, noisy renderings of low-resolution 3DGS coincidentally resemble the natural image statistics favored by no-reference quality metrics, whereas the smoother and more coherent outputs of Mip-Splatting and SR-based methods appear less “natural” under such measures. In contrast, our approach delivers substantially higher perceptual fidelity and cross-view consistency, producing sharper and more realistic high-resolution reconstructions that align closely with both human perception and ground-truth images.
Figure 6: Effect of ratio threshold on Tanks & Temples [Knapitsch2017tandt]. Weight maps, where bright regions indicate higher SR influence, are shown below the corresponding ratio thresholds. $\tau{=}0$ and $\tau{=}\infty$ correspond to zero and full use of super-resolution. SR is initially helpful in improving rendering quality, but excessive use worsens results. The effect of ratio threshold on different scenes is analyzed in Appendix A.2.
Table 2 reports results on Deep Blending [DeepBlending2018] and Mip-NeRF 360 [barron2022mipnerf360]. Our method achieves the strongest performance across all metrics on Deep Blending. On Mip-NeRF 360, it outperforms SRGS on every metric except LPIPS; however, both SR-based approaches are surpassed by Mip-Splatting. This advantage arises from the characteristics of Mip-NeRF 360, where smooth, circular camera trajectories provide dense multi-view coverage and minimal undersampling. In such well-sampled settings, Mip-Splatting’s multi-scale anti-aliasing filters effectively preserve radiance energy and suppress aliasing, whereas SR-based methods may introduce slight inconsistencies when enhancing already well-resolved regions. Additionally, the LR images in Mip-Nerf 360 are approximately twice as large as those in the other datasets. Since most high-frequency details are already preserved at this scale, SR offers minimal improvement. Overall, SplatSuRe achieves the highest performance on Tanks & Temples and Deep Blending and outperforms SRGS on Mip-NeRF 360. Results for $8\times$ SR, additional visualizations, and per-scene metrics for Tables 1 and 2 can be found in Appendix A.1, A.3, and A.4.
@@ 7.1 Ablation on Ratio Threshold
Scenes with different geometries and scales exhibit varying distributions of high and low-frequency content across training views. To account for this variability, we examine the effect of the ratio threshold used to determine where SR is applied. As shown in Figure 6, introducing a small amount of SR initially improves both PSNR and LPIPS by recovering fine details in undersampled regions. However, applying SR too aggressively leads to inconsistencies across views, degrading overall performance. This trend is most pronounced in scenes with varied camera-object distances, where sampling density differs significantly across views. Based on this analysis, we select ratio threshold $\tau{=}1.1$, which provides the best trade-off between sharpness and consistency across scene, in our main experiments. See Appendix A.2 for a detailed analysis on different scenes.
@@ 7.2 Ablation on Super-Resolution Model
We evaluate our method using two single-image super-resolution (SISR) models: SwinIR[liang2021swinir] and StableSR[wang2024exploiting]. SwinIR is optimized for high fidelity under pixel-based metrics such as PSNR, producing accurate yet often over-smoothed reconstructions. In contrast, StableSR incorporates diffusion-based generative priors to synthesize sharper, more detailed, and perceptually realistic images, sometimes at the expense of lower PSNR. As shown in Table 3, our approach improves performance across both models, demonstrating that it is agnostic to the choice of SISR backbone. Notably, our gains are larger with StableSR, as its higher perceptual quality comes at the cost of multi-view inconsistencies. Across both methods, StableSR produces better perceptual quality metrics, motivating our use of StableSR for the main experiments where perceptual realism is prioritized alongside reconstruction fidelity.
Table 3: Comparison of SwinIR [liang2021swinir] and StableSR [wang2024exploiting] on Tanks & Temples [Knapitsch2017tandt]. Experiments are performed at $4\times$ super-resolution using ratio threshold $\tau{=}1.1$. Our method outperforms SRGS with either model. While SwinIR achieves higher PSNR due to its conservative reconstruction, we choose StableSR for our main experiments for its superior perceptual quality.
Method SwinIR [liang2021swinir] StableSR [wang2024exploiting]
@@ 8 Limitations and Future Work
Our method reduces multi-view inconsistencies by suppressing super-resolution in regions already well-captured by low-resolution views. However, this conservative strategy may miss cases where SR could still offer useful refinements. For example, many SR models generate stable, visually sharp detail along high-contrast boundaries. Incorporating adaptive mechanisms that selectively apply SR in such regions could further enhance reconstruction quality. Additionally, our framework currently operates at a single upsampling level, modulating only the strength of SR supervision. Extending this to a multi-scale formulation – where smaller upsampling ratios offer more consistent detail than larger ones – may offer finer control over SR integration and lead to additional gains in sharpness.
@@ A.1 Results for 8× Super-Resolution
Table 4 reports quantitative results at $8\times$ super-resolution for Tanks & Temples [Knapitsch2017tandt]. We downsample the original $1920\times 1080$ images by $16\times$ and upsample by $8\times$ to half the native resolution. Our method achieves the strongest performance across most metrics (SSIM, PSNR, LPIPS, FID and DreamSim). For MUSIQ and NIQE, 3DGS (LR) attains higher scores, consistent with the NIQE behavior discussed in Section 6. Our CMMD score ranks fourth among the compared methods. Prior work [jayasumana2024rethinkingfidbetterevaluation] shows that CMMD is sensitive to high-frequency distortions introduced by noise in the latent embedding space. We hypothesize that the sharpening effects like mild aliasing or overly crisp edges introduced at large upsampling factors may be interpreted by CMMD as distortions, even though they improve perceptual quality and are reflected positively by other metrics. Since no single metric fully characterizes visual fidelity, we report a broad suite of metrics to provide a more comprehensive assessment of our method.
@@
A.2 Additional Ratio Threshold Analysis
@@
Different scenes exhibit distinct behaviors as the ratio threshold and the corresponding amount of super-resolution (SR) information are increased. Most scenes benefit from a moderate amount of SR, but experience a sharp drop in image quality when it is excessively applied. Others plateau in image quality or continue improving, although with diminishing returns. We visualize these trends by plotting PSNR and LPIPS across ratio thresholds for three representative scenes from each category.
@@
Figure 7 illustrates scenes that benefit from an optimal amount of SR. These scenes exhibit a characteristic pattern in which image quality initially improves with moderate SR, but decreases sharply when it is excessively applied. Our method identifies the most poorly sampled regions and selectively applies SR to them, yielding a substantial initial quality boost. However, applying excessive SR introduces multi-view inconsistencies that rapidly degrades image quality. This behavior appears consistently across most scenes, supporting our hypothesis that selectively applying SR is more beneficial than applying it uniformly.
@@
Figure 8 presents scenes that plateau in image quality or continue to benefit slightly as the amount of SR is increased. Our method identifies the most poorly sampled regions and selectively applies SR to them, producing the sharp early quality gains at lower ratio thresholds. Beyond this point, increasing the amount of SR yields diminishing returns or no improvement. This is particularly observed in scenes where the input images already contain substantial high-frequency detail. In such cases, SR produces simpler sharpening or edge-enhancement effects rather than hallucinating new structure, making uniform application less harmful and sometimes marginally beneficial. This behavior is especially common in outdoor scenes in Mip-NeRF 360 [barron2022mipnerf360], where the downsampled $\sim 500\times 375$ images remain relatively high-resolution and therefore do not exhibit the multi-view inconsistencies that typically arise when excessive SR is applied.
A.3 Additional Visualizations
@@
Figures 9 and 10 present additional qualitative results on Tanks & Temples [Knapitsch2017tandt]. Consistent with the examples in Figures 1 and 5 discussed in Section 6, SplatSuRe produces sharper reconstructions than competing methods while reducing artifacts and preserving smoothness in uniformly textured regions.
A.4 Per-Scene Metrics
@@
Tables 6, 7, 8, 9, 10, 11, 12, and 13 report per-scene SSIM, PSNR, LPIPS, FID, CMMD, DreamSim, MUSIQ, and NIQE results on Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360] for all methods evaluated in Section 5. Across individual scenes, we observe the same trends as in the averaged results reported in Tables 1 and 2 in Section 6: SplatSuRe achieves the strongest performance on Tanks & Temples and Deep Blending, and outperforms SRGS on Mip-NeRF 360.
@@
Method Tanks & Temples [Knapitsch2017tandt]
@@
3DGS (LR) [kerbl3Dgaussians] 0.537 16.40 0.473 172.44 3.795 0.2427 46.870 3.976
StableSR [wang2024exploiting] 0.664 21.49 0.406 99.80 1.878 0.1100 38.929 6.293
Mip-Splatting [Yu2024MipSplatting] 0.661 21.53 0.428 109.01 1.941 0.1183 32.417 6.645
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.674 21.97 0.399 92.91 1.891 0.0981 38.322 6.138
Ours + StableSR [wang2024exploiting] 0.692 22.48 0.378 77.17 2.007 0.0774 44.428 5.341
@@
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
@@
3DGS (LR) [kerbl3Dgaussians] 0.783 24.50 0.404 2.051 0.1280 0.509 18.17 0.491 2.395 0.1337
StableSR [wang2024exploiting] 0.821 26.68 0.383 1.159 0.0705 0.630 24.23 0.445 1.096 0.0546
Mip-Splatting [Yu2024MipSplatting] 0.831 27.24 0.391 1.127 0.0689 0.656 24.80 0.420 0.610 0.0243
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.830 27.35 0.377 1.079 0.0622 0.648 24.88 0.422 0.665 0.0320
Ours + StableSR [wang2024exploiting] 0.843 28.03 0.357 0.879 0.0489 0.655 25.08 0.406 0.546 0.0231
(a) ballroom – Tanks & Temples [Knapitsch2017tandt]
(b) kitchen – Mip-NeRF 360 [barron2022mipnerf360]
(c) francis – Tanks & Temples [Knapitsch2017tandt]
@@
Figure 7: Representative scenes that benefit from an optimal amount of super-resolution. Top: Image quality vs. ratio threshold plots. Bottom: ground truth images illustrating scene structure for (a) ballroom from Tanks & Temples [Knapitsch2017tandt], (b) kitchen from Mip-NeRF 360 [barron2022mipnerf360] and (c) francis from Tanks & Temples. Applying SR to the most poorly sampled regions yields large gains in image quality, whereas excessive SR introduces multi-view inconsistencies that sharply degrade quality. Most scenes exhibit this behavior, supporting our hypothesis that selectively applying SR is more beneficial than applying it uniformly.
(a) bicycle – Mip-NeRF 360 [barron2022mipnerf360]
(b) garden – Mip-NeRF 360 [barron2022mipnerf360]
(c) stump – Mip-NeRF 360 [barron2022mipnerf360]
@@
Figure 8: Representative scenes that plateau in image quality or continue to benefit from increased amounts of super-resolution. Top: Image quality vs. ratio threshold plots. Bottom: ground truth images illustrating scene structure for (a) bicycle, (b) garden, and (c) stump from Mip-NeRF 360 [barron2022mipnerf360]. Applying SR to the most poorly sampled regions yields large gains in image quality, while further increasing SR provides diminishing returns or no improvement. This is particularly observed in scenes where the input images already contain substantial high-frequency detail. In such cases, SR produces simpler sharpening or edge-enhancement effects rather than hallucinating new structure, making uniform application less harmful and sometimes marginally beneficial.
@@
Figure 9: Additional qualitative results on Tanks & Temples [Knapitsch2017tandt]. Experiments are performed at $4\times$ super-resolution with ratio threshold $\tau{=}1.1$. Compared to Mip-Splatting [Yu2024MipSplatting] and SRGS [feng2024srgssuperresolution3dgaussian], our method produces sharper, more faithful reconstructions that better align with ground truth while maintaining cross-view consistency. It preserves high-frequency patterns (orange boxes on statue), fine details in text (purple box on vehicle), and reduces artifacts while preserving sharpness (green boxes on curtains, blue boxes in playground).
@@
Figure 10: Additional qualitative results on Tanks & Temples [Knapitsch2017tandt]. Experiments are performed at $4\times$ super-resolution with ratio threshold $\tau{=}1.1$. Compared to Mip-Splatting [Yu2024MipSplatting] and SRGS [feng2024srgssuperresolution3dgaussian], our method produces sharper, more faithful reconstructions that better align with ground truth while maintaining cross-view consistency. It preserves high-frequency geometry (pink boxes on barn) and retains sharp texture (yellow boxes on lighthouse, red boxes on horse statue, blue boxes on tank).
@@
Table 6: SSIM $\uparrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360]. Experiments are performed at $4\times$ super-resolution using ratio threshold $\tau{=}1.1$. The best, second best and third best entries are highlighted.
Method Tanks & Temples [Knapitsch2017tandt]
@@
3DGS (LR) [kerbl3Dgaussians] 0.807 0.527 0.666 0.578 0.752 0.729 0.742 0.628 0.647 0.735 0.715 0.602 0.566 0.648 0.779 0.649 0.682 0.675 0.579
StableSR [wang2024exploiting] 0.849 0.646 0.709 0.702 0.819 0.771 0.777 0.720 0.725 0.810 0.761 0.686 0.753 0.705 0.845 0.759 0.744 0.811 0.683
Mip-Splatting [Yu2024MipSplatting] 0.838 0.681 0.687 0.729 0.847 0.790 0.789 0.746 0.730 0.836 0.762 0.704 0.788 0.733 0.855 0.786 0.753 0.827 0.689
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.861 0.671 0.713 0.734 0.846 0.794 0.789 0.742 0.746 0.835 0.773 0.703 0.780 0.729 0.857 0.784 0.761 0.829 0.704
Ours + StableSR [wang2024exploiting] 0.859 0.694 0.717 0.746 0.859 0.808 0.807 0.764 0.752 0.846 0.786 0.714 0.795 0.743 0.866 0.798 0.781 0.835 0.716
Method Deep Blending [DeepBlending2018] Mip-NeRF 360 [barron2022mipnerf360]
@@
3DGS (LR) [kerbl3Dgaussians] 0.828 0.843 0.527 0.776 0.772 0.429 0.559 0.719 0.836 0.581 0.516
StableSR [wang2024exploiting] 0.837 0.854 0.604 0.848 0.833 0.497 0.654 0.751 0.858 0.669 0.571
Mip-Splatting [Yu2024MipSplatting] 0.855 0.874 0.676 0.917 0.872 0.544 0.719 0.870 0.897 0.740 0.596
SRGS [feng2024srgssuperresolution3dgaussian] + StableSR [wang2024exploiting] 0.851 0.870 0.653 0.896 0.863 0.533 0.692 0.825 0.888 0.671 0.590
Ours + StableSR [wang2024exploiting] 0.867 0.877 0.646 0.906 0.867 0.503 0.690 0.869 0.896 0.706 0.582
Table 7: PSNR $\uparrow$ on each scene in Tanks & Temples [Knapitsch2017tandt], Deep Blending [DeepBlending2018], and Mip-NeRF 360 [barron2022mipnerf360].
Method Tanks & Temples [Knapitsch2017tandt]

### v2

@@ 6 Results
Figures 1 and 5 present representative renderings from our method. Our approach produces consistently sharper reconstructions while preserving smoothness in uniformly textured regions. In contrast, Mip-Splatting yields overly blurred results because it is trained only on LR inputs and its 3D anti-aliasing filter excessively blurs all Gaussians. SRGS recovers higher-frequency content in some areas but remains constrained by the underlying SR images – regions where the SR output is consistent appear sharp, whereas areas with view-inconsistent SR predictions become noticeably blurred or distorted.
@@
Table 1 presents results on Tanks & Temples [15]. Our method achieves the highest image quality across nearly all metrics, demonstrating both its effectiveness and robustness. Only 3DGS attains a better NIQE score. Similarly, SRGS ranks second across most reference and perceptual metrics but lags behind 3DGS in MUSIQ and NIQE. This trend arises because the aliased, noisy renderings of low-resolution 3DGS coincidentally resemble the natural image statistics favored by no-reference quality metrics, whereas the smoother and more coherent outputs of Mip-Splatting and SR-based methods appear less “natural” under such measures. In contrast, our approach delivers substantially higher perceptual fidelity and cross-view consistency, producing sharper and more realistic high-resolution reconstructions that align closely with both human perception and ground-truth images.
Table 2 reports results on Deep Blending [8] and Mip-NeRF 360 [1]. Our method achieves the strongest performance across all metrics on Deep Blending. On Mip-NeRF 360, it outperforms SRGS on every metric except LPIPS; however, both SR-based approaches are surpassed by Mip-Splatting. This arises from the characteristics of Mip-NeRF 360, where smooth, circular camera trajectories provide dense multi-view coverage and minimal undersampling. In these well-sampled settings, Mip-Splatting’s multi-scale anti-aliasing filters effectively preserve radiance energy and suppress aliasing, whereas SR-based methods may introduce slight inconsistencies when enhancing already well-resolved regions. Additionally, the LR images in Mip-NeRF 360 are roughly twice as large as those in the other datasets, preserving most high-frequency details and leaving little room for SR improvement. Overall, SplatSuRe achieves the highest performance on Tanks & Temples and Deep Blending and outperforms SRGS on Mip-NeRF 360. Results for $8\times$ SR, additional visualizations, and per-scene metrics for Tables 1 and 2 can be found in Appendix A.1, A.4, and A.5. Appendix A.2 presents a unified training pipeline that merges the LR initialization and SR refinement stages to achieve comparable performance with the same training budget as single-stage baselines.
Figure 6: Effect of ratio threshold on Tanks & Temples [15]. Weight maps, where bright regions indicate higher SR influence, are shown below the corresponding ratio thresholds. $\tau{=}0$ and $\tau{=}\infty$ correspond to zero and full use of super-resolution. SR is initially helpful in improving rendering quality, but excessive use worsens results. The effect of ratio threshold on different scenes is analyzed in Appendix A.3.
@@ 7.1 Ablation on Ratio Threshold
Scenes with different geometries and scales exhibit varying distributions of high and low-frequency content across training views. To account for this variability, we examine the effect of the ratio threshold used to determine where SR is applied. As shown in Figure 6, introducing a small amount of SR initially improves both PSNR and LPIPS by recovering fine details in undersampled regions. However, applying SR too aggressively leads to inconsistencies across views, degrading overall performance. This trend is most pronounced in scenes with varied camera-object distances, where sampling density differs significantly across views. Based on this analysis, we select ratio threshold $\tau{=}1.1$, which provides the best trade-off between sharpness and consistency across scenes, in our main experiments. See Appendix A.3 for a detailed analysis of different scenes.
@@ 7.2 Ablation on Super-Resolution Model
We evaluate our method using two single-image super-resolution (SISR) models: SwinIR [16] and StableSR [27]. SwinIR is optimized for high fidelity under pixel-based metrics such as PSNR, producing accurate yet often over-smoothed reconstructions. In contrast, StableSR incorporates diffusion-based generative priors to synthesize sharper, more detailed, and perceptually realistic images, sometimes at the expense of lower PSNR. As shown in Table 3, our approach improves performance across both models, demonstrating that it is agnostic to the choice of SISR backbone. Notably, our gains are larger with StableSR, as its higher perceptual quality comes at the cost of multi-view inconsistencies. Across both methods, StableSR produces better perceptual quality metrics, motivating our use of StableSR for the main experiments where perceptual realism is prioritized alongside reconstruction fidelity.
Table 3: Comparison of SwinIR [16] and StableSR [27] on Tanks & Temples [15]. Experiments are performed at $4\times$ super-resolution using ratio threshold $\tau{=}1.1$. Our method outperforms SRGS with either model. While SwinIR achieves higher PSNR due to its conservative reconstruction, we choose StableSR for our main experiments for its superior perceptual quality.
Method SwinIR [16] StableSR [27]
@@ 8 Limitations and Future Work
Our method reduces multi-view inconsistencies by suppressing SR in regions already well-captured by LR views, but this conservative strategy may miss useful SR refinements, such as stable detail along high-contrast boundaries. Selectively applying SR in these regions could further improve reconstruction quality. In addition, our framework operates at a single upsampling level. Extending it to a multi-scale formulation could provide finer control over SR integration and improve sharpness. Finally, while SplatSuRe produces multi-view consistent 3D models faithful to ground truth, it could further benefit from advances in SR that reduce multi-view inconsistency in generative outputs.
@@ A.1 Results for 8× Super-Resolution
Table 4 reports quantitative results at $8\times$ super-resolution (SR) for Tanks & Temples [15]. We downsample the original $1920\times 1080$ images by $16\times$ and upsample by $8\times$ to half the native resolution. Our method achieves the best SSIM, PSNR, LPIPS, FID, and DreamSim scores. 3DGS (LR) attains higher MUSIQ and NIQE scores, consistent with the behavior discussed in Section 6. Our CMMD score ranks fourth among the compared methods. Prior work [12] shows that CMMD is sensitive to high-frequency distortions introduced by noise in the embedding space. We hypothesize that sharpening effects like mild aliasing or overly crisp edges at high upsampling factors may be interpreted as distortions by CMMD, even when they improve perceptual quality and are reflected positively by other metrics. As no single metric fully characterizes visual fidelity, we report a broad suite of metrics for a more complete evaluation.
@@
A.2 Unified Training Pipeline
Table 6 evaluates a unified training pipeline that merges the LR initialization and SR refinement stages to avoid the training time overhead introduced by the two-stage pipeline in Section 4. The model is trained with LR images for the first 5K iterations to obtain stable geometry, after which the Gaussian fidelity scores and SR weight maps are computed and training continues with SR supervision for the remaining 25K iterations. The total budget of 30K iterations matches that used by baseline methods using a single-stage pipeline. This unified approach achieves performance comparable to the original two-stage formulation reported in Tables 1 and 2, indicating that our SplatSuRe objective remains effective as a single continuous training schedule.
A.3 Additional Ratio Threshold Analysis
@@
Different scenes exhibit distinct behaviors as the ratio threshold and corresponding amount of super-resolution (SR) information increase. Most scenes benefit from a moderate amount of SR but experience a sharp drop in image quality when it is excessively applied, while others plateau or continue improving with diminishing returns. We visualize these trends by plotting PSNR and LPIPS across ratio thresholds for three representative scenes in each category.
@@
Figure 7 illustrates scenes that benefit from an optimal amount of SR. Image quality initially improves with moderate SR but decreases sharply when it is excessively applied. Our method identifies the most poorly sampled regions and selectively applies SR to them, yielding a substantial initial quality boost. However, applying excessive SR introduces multi-view inconsistencies that rapidly degrade image quality. This behavior appears consistently across most scenes, supporting our hypothesis that selectively applying SR is more beneficial than applying it uniformly.
@@
Figure 8 presents scenes that plateau in image quality or continue improving slightly as the amount of SR is increased. Our selective SR method produces sharp early quality gains at lower ratio thresholds, after which applying additional SR yields diminishing returns or no improvement. In particular, this occurs in scenes where the input images already contain substantial high-frequency detail and SR produces simpler sharpening or edge enhancement effects rather than hallucinating new structure, making uniform application less harmful and sometimes marginally beneficial. This behavior is especially common in outdoor scenes in Mip-NeRF 360 [1], where the downsampled $\sim 500\times 375$ images remain relatively high-resolution and therefore do not exhibit the multi-view inconsistencies that typically arise when excessive SR is applied.
A.4 Additional Visualizations
@@
Figures 9 and 10 present additional qualitative results on Tanks & Temples [15]. Consistent with the examples in Figures 1 and 5 discussed in Section 6, SplatSuRe produces sharper reconstructions than competing methods while reducing artifacts and preserving smoothness in uniformly textured regions.
A.5 Per-Scene Metrics
@@
Tables 7, 8, 9, 10, 11, 12, 13, and 14 report per-scene SSIM, PSNR, LPIPS, FID, CMMD, DreamSim, MUSIQ, and NIQE results on Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1] for all methods evaluated in Section 5. Across individual scenes, we observe the same trends as in the averaged results presented by Tables 1 and 2 in Section 6: SplatSuRe achieves the strongest performance on Tanks & Temples and Deep Blending and outperforms SRGS on Mip-NeRF 360.
@@
Method Tanks & Temples [15]
@@
3DGS (LR) [14] 0.537 16.40 0.473 172.44 3.795 0.2427 46.870 3.976
3DGS [14] + StableSR [27] 0.664 21.49 0.406 99.80 1.878 0.1100 38.929 6.293
Mip-Splatting [33] 0.661 21.53 0.428 109.01 1.941 0.1183 32.417 6.645
SRGS [6] + StableSR [27] 0.674 21.97 0.399 92.91 1.891 0.0981 38.322 6.138
Ours + StableSR [27] 0.692 22.48 0.378 77.17 2.007 0.0774 44.428 5.341
@@
Method Deep Blending [8] Mip-NeRF 360 [1]
@@
3DGS (LR) [14] 0.783 24.50 0.404 2.051 0.1280 0.509 18.17 0.491 2.395 0.1337
3DGS [14] + StableSR [27] 0.821 26.68 0.383 1.159 0.0705 0.630 24.23 0.445 1.096 0.0546
Mip-Splatting [33] 0.831 27.24 0.391 1.127 0.0689 0.656 24.80 0.420 0.610 0.0243
SRGS [6] + StableSR [27] 0.830 27.35 0.377 1.079 0.0622 0.648 24.88 0.422 0.665 0.0320
Ours + StableSR [27] 0.843 28.03 0.357 0.879 0.0489 0.655 25.08 0.406 0.546 0.0231
Table 6: Quantitative comparison of our unified and two-stage pipelines at $\mathbf{4\times}$ SR across Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1]. Experiments are performed using ratio threshold $\tau=1.1$. The best entry is bolded. The unified pipeline achieves similar performance to the two-stage approach while requiring less training time.
Dataset Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ CMMD $\downarrow$ DreamSim $\downarrow$ MUSIQ $\uparrow$ NIQE $\downarrow$
Tanks & Temples [15] Two-Stage 0.784 23.81 0.272 37.72 1.040 0.0413 58.332 3.928
Unified 0.781 23.70 0.271 36.02 1.006 0.0400 59.701 3.705
Deep Blending [8] Two-Stage 0.872 29.01 0.306 44.14 0.496 0.0330 53.019 5.411
Unified 0.871 28.96 0.304 41.70 0.483 0.0326 52.983 5.286
Mip-NeRF 360 [1] Two-Stage 0.740 26.34 0.323 27.56 0.339 0.0179 54.366 3.934
Unified 0.758 26.69 0.304 26.18 0.271 0.0164 56.773 3.715
(a) ballroom – Tanks & Temples [15]
(b) kitchen – Mip-NeRF 360 [1]
(c) francis – Tanks & Temples [15]
@@
Figure 7: Representative scenes that benefit from an optimal amount of super-resolution. Top: Image quality vs. ratio threshold plots. Bottom: ground truth images illustrating scene structure for (a) ballroom from Tanks & Temples [15], (b) kitchen from Mip-NeRF 360 [1] and (c) francis from Tanks & Temples. Applying SR to the most poorly sampled regions yields large gains in image quality, whereas excessive SR introduces multi-view inconsistencies that sharply degrade quality. Most scenes exhibit this behavior, supporting our hypothesis that selectively applying SR is more beneficial than applying it uniformly.
(a) bicycle – Mip-NeRF 360 [1]
(b) garden – Mip-NeRF 360 [1]
(c) stump – Mip-NeRF 360 [1]
@@
Figure 8: Representative scenes that plateau in image quality or continue to benefit from increased amounts of super-resolution. Top: Image quality vs. ratio threshold plots. Bottom: ground truth images illustrating scene structure for (a) bicycle, (b) garden, and (c) stump from Mip-NeRF 360 [1]. Applying SR to the most poorly sampled regions yields large gains in image quality, while further increasing SR yields diminishing returns or no improvement. In particular, this occurs in scenes where the input images already contain substantial high-frequency detail and SR produces simpler sharpening or edge-enhancement effects rather than hallucinating new structure, making uniform application less harmful and sometimes marginally beneficial.
@@
Figure 9: Additional qualitative results on Tanks & Temples [15]. Experiments are performed at $4\times$ super-resolution with ratio threshold $\tau{=}1.1$. Compared to Mip-Splatting [33] and SRGS [6], our method produces sharper, more faithful reconstructions that better align with ground truth while maintaining cross-view consistency. It preserves high-frequency patterns (orange boxes on statue), fine details in text (purple box on vehicle), and reduces artifacts while preserving sharpness (green boxes on curtains, blue boxes in playground).
@@
Figure 10: Additional qualitative results on Tanks & Temples [15]. Experiments are performed at $4\times$ super-resolution with ratio threshold $\tau{=}1.1$. Compared to Mip-Splatting [33] and SRGS [6], our method produces sharper, more faithful reconstructions that better align with the ground truth while maintaining cross-view consistency. It preserves high-frequency geometry (pink boxes on barn) and retains sharp texture (yellow boxes on lighthouse, red boxes on horse statue, blue boxes on tank).
@@
Table 7: SSIM $\uparrow$ on each scene in Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1]. Experiments are performed at $4\times$ super-resolution using ratio threshold $\tau{=}1.1$. The best, second best and third best entries are highlighted.
Method Tanks & Temples [15]
@@
3DGS (LR) [14] 0.807 0.527 0.666 0.578 0.752 0.729 0.742 0.628 0.647 0.735 0.715 0.602 0.566 0.648 0.779 0.649 0.682 0.675 0.579
3DGS [14] + StableSR [27] 0.849 0.646 0.709 0.702 0.819 0.771 0.777 0.720 0.725 0.810 0.761 0.686 0.753 0.705 0.845 0.759 0.744 0.811 0.683
Mip-Splatting [33] 0.838 0.681 0.687 0.729 0.847 0.790 0.789 0.746 0.730 0.836 0.762 0.704 0.788 0.733 0.855 0.786 0.753 0.827 0.689
SRGS [6] + StableSR [27] 0.861 0.671 0.713 0.734 0.846 0.794 0.789 0.742 0.746 0.835 0.773 0.703 0.780 0.729 0.857 0.784 0.761 0.829 0.704
Ours + StableSR [27] 0.859 0.694 0.717 0.746 0.859 0.808 0.807 0.764 0.752 0.846 0.786 0.714 0.795 0.743 0.866 0.798 0.781 0.835 0.716
Method Deep Blending [8] Mip-NeRF 360 [1]
@@
3DGS (LR) [14] 0.828 0.843 0.527 0.776 0.772 0.429 0.559 0.719 0.836 0.581 0.516
3DGS [14] + StableSR [27] 0.837 0.854 0.604 0.848 0.833 0.497 0.654 0.751 0.858 0.669 0.571
Mip-Splatting [33] 0.855 0.874 0.676 0.917 0.872 0.544 0.719 0.870 0.897 0.740 0.596
SRGS [6] + StableSR [27] 0.851 0.870 0.653 0.896 0.863 0.533 0.692 0.825 0.888 0.671 0.590
Ours + StableSR [27] 0.867 0.877 0.646 0.906 0.867 0.503 0.690 0.869 0.896 0.706 0.582
Table 8: PSNR $\uparrow$ on each scene in Tanks & Temples [15], Deep Blending [8], and Mip-NeRF 360 [1].
Method Tanks & Temples [15]